# KOALA++: Efficient Kalman-Based Optimization of Neural Networks with Gradient-Covariance Products

### v1

@@ Abstract
We propose KOALA++, a scalable Kalman-based optimization algorithm that explicitly models structured gradient uncertainty in neural network training. Unlike second-order methods, which rely on expensive second order gradient calculation, our method directly estimates the parameter covariance matrix by recursively updating compact gradient covariance products. This design improves upon the original KOALA framework that assumed diagonal covariance by implicitly capturing richer uncertainty structure without storing the full covariance matrix and avoiding large matrix inversions. Across diverse tasks, including image classification and language modeling, KOALA++ achieves accuracy on par or better than state-of-the-art first- and second-order optimizers while maintaining the efficiency of first-order methods.
@@ 1 Introduction
To mitigate these issues, second-order optimizers such as K-FAC (martens2015optimizing, ) and Shampoo (gupta2018shampoo, ) introduce curvature-aware updates by approximating the Hessian or Fisher information matrix. These methods often improve convergence and generalization but require costly matrix operations, limiting their scalability. Recently, AdaFisher (gomes2024adafisher, ) has emerged as a more scalable alternative, leveraging block-diagonal Kronecker approximations of the Fisher information matrix to reduce overhead while preserving second-order benefits. Despite its effectiveness, AdaFisher fundamentally depends on the quality of the Fisher matrix estimation, which remains sensitive to noise and mini-batch variability.
@@
An alternative and theoretically grounded direction stems from Kalman filtering. In Bayesian estimation theory, the posterior covariance matrix of an unbiased estimator defines the tightest achievable lower bound on parameter uncertainty, as formalized by the Cramér–Rao bound (rao1992information, ; cramer1999mathematical, ). While the Fisher information matrix serves as an inverse bound proxy, it does not directly capture the true uncertainty in parameters. Kalman filters, on the other hand, propagate the full covariance matrix of the estimate through time, providing a more faithful representation of uncertainty – particularly in noisy or nonstationary environments.
@@
Motivated by this insight, KOALA (davtyan2022koala, ) proposed to frame stochastic optimization as a Kalman filtering problem. It models parameter updates as recursive state estimation, using a covariance matrix to adjust step sizes adaptively. However, KOALA’s original formulation relies on overly simplistic approximations, such as diagonal covariances and idealized gradient assumptions, which limit its practical effectiveness.
@@
In this paper, we introduce KOALA++, a principled extension of the KOALA framework that explicitly maintains structured gradient covariance information. Rather than approximating second-order curvature via Fisher matrices, KOALA++ tracks the evolution of uncertainty directly through a recursive update at each training iteration $k$ of the gradient-covariance product $v_{k}=H_{k}P_{k-1}$, where $H_{k}$ is the gradient of the loss function and $P_{k-1}$ is the prior estimate of the parameter covariance. By doing so, KOALA++ effectively avoids expensive matrix inversions and storing the full covariance matrix $P_{k-1}$, while benefiting from the lifted constraints on its structure. Moreover, KOALA++ retains scalability akin to first-order methods, while achieving optimization dynamics and final performance competitive with state-of-the-art second-order methods.
@@
Our main contributions are summarized as follows: We propose KOALA++, a Kalman-inspired optimizer that explicitly propagates structured gradient uncertainty via covariance modeling. We derive an efficient recursive update for the surrogate $v_{k}$ based on a symmetric minimum-norm solution to the covariance estimation problem. We implement KOALA++ with minimal overhead and validate its effectiveness across diverse architectures and tasks, including ResNet family of models he2016deep , vision transformers liu2021swin , and language models. Extensive experiments show that KOALA++ consistently matches or exceeds the accuracy of first- and second-order baselines while offering improved stability and a better efficiency – performance trade-off.
@@ 2.1 Optimization Methods: A Brief Overview
First-order optimizers such as SGD robbins1951stochastic and Adam kingma2015adam are widely adopted for their scalability and simplicity, but they often ignore the structure of gradient noise and require manually tuned learning rate schedules.
@@
Second-order optimization methods, including K-FAC martens2015optimizing , Shampoo gupta2018shampoo , and more recently AdaFisher gomes2024adafisher , seek to address these shortcomings by explicitly incorporating curvature information. Nevertheless, practical implementations rely heavily on fixed approximations, such as block-diagonal or Kronecker-factored representations of the Hessian or Fisher matrices, to maintain computational feasibility. Such static approximations fail to adapt dynamically to the non-stationary curvature and noise structure inherent in deep network training.
2.2 Kalman Filtering Perspective for Optimization
To overcome these limitations, recent work proposes modeling optimization as a probabilistic filtering process, inspired by Kalman filtering kalman1960new . In fact, the idea of applying Kalman filtering to neural network optimization is not new. Several earlier works have explored different ways of integrating filtering theory into learning systems. For example, Ismail et al. ismail2018estimation propose a scalar Kalman update rule for stochastic optimization; Shashua and Mannor shashua2019trust apply Kalman filtering in reinforcement learning for value approximation; and Ollivier ollivier2019extended connects the Extended Kalman Filter (EKF) to online natural gradient methods via Amari’s framework. Moreover, de Vilmarest and Wintenberger de2021stochastic apply EKF to logistic regression, and Takedo et al. takenga2004comparison compare GD with Kalman-style variants in control settings.
While these approaches illustrate the versatility of Kalman-based ideas, most are either narrowly scoped to specific domains (e.g., reinforcement learning, linear models), or suffer from limited scalability due to full covariance computation. Among these, KOALA davtyan2022koala stands out as a practical and generalizable solution for deep learning. By interpreting gradient descent as a scalar state estimation problem, KOALA introduces tractable variants that scale to large neural networks using diagonal or isotropic covariance approximations. For a more direct comparison with our method, we report the complete descriptions of KOALA-V and KOALA-M variants in Appendix A.
@@ 2.3 Motivation for KOALA++: Towards Structured Covariance Updates
However, the simplifications made in KOALA also constrain their ability to model structured noise or adapt to anisotropic curvature. In particular, KOALA assumes $P_{k}\approx\sigma^{2}I$ and approximates $H_{k}^{\top}H_{k}$ using gradient norms, which may be suboptimal in noisy or high-dimensional settings.
Our proposed extension, KOALA++, addresses these limitations by explicitly tracking a directional covariance projection $v_{k}:=H_{k}P_{k-1}$ and updating it via a recursive formulation of least squares. This approach maintains the efficiency of first-order updates while offering richer uncertainty modeling than prior filtering-based optimizers.
@@ 3.1 Problem Formulation and Kalman Filtering Notation
The training of neural networks is fundamentally an optimization problem, where given a dataset ${\cal D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, the objective of the learning task is to adjust the model’s parameters $\theta$ such that the network’s predictions $f(x_{i};\theta)$ minimize the empirical risk
$\min_{\theta}{\cal L}(\theta),\quad{\cal L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}l(f(x_{i};\theta),y_{i}),$ (1)
where $l$ is a chosen loss function. However, in modern deep learning applications, $N$ is often large, making full-batch optimization computationally infeasible. As a result, practitioners rely on stochastic optimization methods that operate on minibatches of data, using estimates of the loss
@@
where $B_{k}$ is a set of indices in the $k$-th minibatch and $m=|B_{k}|$ is the minibatch size. While minibatch-based approaches such as SGD robbins1951stochastic and its adaptive variants are effective and scalable, they introduce noise due to the stochastic nature of loss and gradient estimates.
@@
$S_{k}=H_{k}(P_{k-1}+QI)H_{k}^{\top}+R$
$K_{k}=(P_{k-1}+QI)H_{k}^{\top}S_{K}^{-1}$
$P_{k}=(I-K_{k}H_{k})(P_{k-1}+QI)$
$\theta_{k}=\theta_{k-1}+K_{k}(L_{k}^{\text{target}}-L_{k}(\theta_{k-1}))$
To mitigate this issue, a recent approach, KOALA davtyan2022koala , casts neural network training as a recursive state estimation problem using Kalman filtering kalman1960new , a principled Bayesian framework for recursively estimating the hidden state of a system from noisy observations. At each iteration $k$, it combines prior estimates with new measurements to update the state $\theta_{k}$ and its associated uncertainty in the form of a covariance matrix $P_{k}$. Importantly, by keeping track of an estimate of the system’s state’s covariance, Kalman filtering is essentially implicitly using second order information about the state variables. The standard Extended Kalman Filter (EKF) generalizes this to non-linear systems using local linearization. In KOALA, the model parameters $\theta_{k}$ are treated as latent states with stationary evolution under process noise, and the minibatch loss function $L_{k}$ serves as a noisy observation of a target loss value $L_{k}^{\text{target}}$. Under this settings, the EKF equations can simplified as in Table 1. To remain tractable in high-dimensional settings, KOALA simplifies the EKF update by:
Assuming an identity-scaled covariance matrix $P_{k}\approx\sigma_{k}^{2}I$; Approximating $H_{k}^{\top}H_{k}$ with $||H_{k}||^{2}I$; Reducing all matrix operations to scalar computations.
@@
While KOALA demonstrates promising stability and simplicity, its reliance on scalar approximations limits its capacity to capture directional or structured noise in high-dimensional loss landscapes. To address this, we propose KOALA++, which enhances the original framework with an efficient estimation of directional state covariance while preserving the Kalman-based update structure.
@@ 3.2 Kalman Update with $v_{k}$-Based Reparameterization
To overcome the limitations of isotropic covariance assumptions in KOALA, KOALA++ introduces a more expressive formulation that captures directional uncertainty without incurring the full cost of matrix-valued updates. Instead of approximating the full posterior covariance $P_{k}\in\mathbb{R}^{n\times n}$, we propagate a low-rank surrogate
@@
which represents the projection of the parameter covariance along the current measurement direction. This allows KOALA++ to maintain a richer understanding of uncertainty and curvature structure, while keeping the computational complexity comparable to first-order methods.
@@
Notice, that as suggested in davtyan2022koala , the learning rate of the update can be controlled by selecting a specific $L_{k}^{\text{target}}$. For instance, $L_{k}^{\text{target}}=(1-\eta_{k})L_{k}(\theta_{k-1})$ ensures that the update is scaled by $\eta_{k}$.
@@
$=(H_{k}-H_{k}\frac{v_{k-1}^{T}+QH_{k-1}^{T}}{S_{k-1}}H_{k-1})(P_{k-2}+Q)$
$=H_{k}P_{k-2}+H_{k}Q-H_{k}\frac{v_{k-1}^{T}+QH_{k-1}^{T}}{S_{k-1}}v_{k-1}-H_{k}\frac{v_{k-1}^{T}+QH_{k-1}^{T}}{S_{k-1}}H_{k-1}Q.$ (7)
For simplicity, we define
$\lambda_{k}$ $=H_{k}\frac{v_{k-1}^{T}+QH_{k-1}^{T}}{S_{k-1}},$ (8)
which allows to rewrite the update for $v_{k}$ in a more compact way:
@@
In the last equation, all the terms except for the first one can be efficiently calculated at iteration $k$. The first term instead cannot be easily evaluated due to the $P_{k-2}$ term. The objective is to never store it explicitly to avoid the quadratic storage demand. In principle, one can approximate $H_{k}$ with $H_{k-1}$ and replace the problematic term with $v_{k-1}$. However, we observed experimentally that this does not work well due to the noise in the gradients.
To proceed, we aim to retrospectively approximate the prior covariance matrix $P_{k-2}$ instead. We do this by leveraging the relation $v_{k-1}=H_{k-1}P_{k-2}$. However, since this equation is generally underdetermined in high-dimensional settings, we formulate a regularized inverse problem to obtain a unique and tractable solution. Specifically, we adopt the minimum Frobenius-norm formulation
@@
For computational and memory efficiency, we use only one linear constraint, $v_{k-1}=H_{k-1}P_{k-2}$, although more could be used in principle. However, introducing more constraints, e.g. projections of other past gradients $H_{k-2},H_{k-3}$ etc., would also require to memorize more gradient terms and using more calculations to update $v_{k}$. Finally, the use of the minimum Frobenius-norm criterion is to ensure that $P_{k-2}$ has a bounded norm and thus a more stable update for $v_{k}$.
In the next subsection we discuss two variants of our algorithm, depending on the nature of the solution to the optimization problem (10).

### v2

# KOALA++: Efficient Kalman-Based Optimization with Gradient-Covariance Products

@@ Abstract
We propose KOALA++, a scalable Kalman-based optimization algorithm that explicitly models structured gradient uncertainty in neural network training. Unlike second-order methods, which rely on expensive second order gradient calculation, our method directly estimates the parameter covariance matrix by recursively updating compact gradient covariance products. This design improves upon the original KOALA framework that assumed diagonal covariance by implicitly capturing richer uncertainty structure without storing the full covariance matrix and avoiding large matrix inversions. Across diverse tasks, including image classification and language modeling, KOALA++ achieves accuracy on par or better than state-of-the-art first- and second-order optimizers while maintaining the efficiency of first-order methods. The code is publicly available at https://github.com/Sumxiaa/KOALA_Plus_Plus.
@@ 1 Introduction
To mitigate these issues, second-order optimizers such as K-FAC (martens2015optimizing, ) and Shampoo (gupta2018shampoo, ) introduce curvature-aware updates by approximating the Hessian or Fisher information matrix. These methods often improve convergence and generalization, but require costly matrix operations, limiting their scalability. Recently, AdaFisher (gomes2024adafisher, ) has emerged as a more scalable alternative, using block-diagonal Kronecker approximations of the Fisher information matrix to reduce overhead while preserving second-order benefits. Despite its effectiveness, AdaFisher fundamentally depends on the quality of the Fisher matrix estimation, which remains sensitive to noise and mini-batch variability.
@@
An alternative and theoretically grounded direction comes from Kalman filtering. In Bayesian estimation theory, the posterior covariance matrix of an unbiased estimator defines the tightest achievable lower bound on parameter uncertainty, as formalized by the Cramér–Rao bound (rao1992information, ; cramer1999mathematical, ). Although the Fisher information matrix serves as an inverse bound proxy, it does not directly capture the true uncertainty in the parameters. Kalman filters, on the other hand, propagate the full covariance matrix of the estimate through time, providing a more faithful representation of uncertainty – particularly in noisy or nonstationary environments.
@@
Motivated by this insight, KOALA (davtyan2022koala, ) proposed to frame stochastic optimization as a Kalman filtering problem, which models parameter updates as recursive state estimation, using a covariance matrix to adjust step sizes adaptively. However, KOALA’s original formulation relies on overly simplistic approximations, such as diagonal covariances and idealized gradient assumptions, which limit its practical effectiveness.
@@
In this paper, we introduce KOALA++, a computationally efficient extension of the KOALA framework that explicitly maintains structured gradient covariance information. KOALA++ tracks the evolution of uncertainty directly through a recursive update at each training iteration $k$ of the gradient-covariance product $v_{k}=H_{k}P_{k-1}$, which we call surrogate, where $H_{k}$ is the gradient of the loss function and $P_{k-1}$ is the prior estimate of the parameter covariance. The key idea in our method is to approximate the recursive gradient-covariance update so that it never uses the covariance matrix $P_{k-1}$. By doing so, KOALA++ effectively avoids expensive matrix inversions and storing the entire covariance matrix $P_{k-1}$, while benefiting from the lifted constraints on its structure. Moreover, KOALA++ retains scalability akin to first-order methods, while achieving optimization dynamics and final performance competitive with state-of-the-art second-order methods.
@@
Our main contributions are summarized as follows: We propose KOALA++, a Kalman-inspired optimizer that explicitly propagates structured gradient uncertainty via covariance modeling. We derive an efficient recursive update for the surrogate $v_{k}$ based on a minimum-norm solution to the covariance estimation problem. We implement KOALA++ with minimal overhead and validate its effectiveness across diverse architectures and tasks, including the ResNet family he2016deep , vision transformers liu2021swin , and language models. Extensive experiments show that KOALA++ consistently matches or exceeds the accuracy of first- and second-order baselines while offering improved stability and a better efficiency/performance trade-off.
@@ 2.1 Optimization Methods: A Brief Overview
First-order optimizers such as SGD robbins1951stochastic and Adam kingma2015adam are widely adopted for their scalability and simplicity, but often ignore the structure of gradient noise and require manually tuned learning rate schedules.
@@
Second-order optimization methods, including K-FAC martens2015optimizing , Shampoo gupta2018shampoo , and more recently AdaFisher gomes2024adafisher , seek to address these shortcomings by explicitly incorporating curvature information. Nevertheless, practical implementations rely heavily on fixed approximations, such as block-diagonal or Kronecker-factored representations of the Hessian or Fisher matrices, to maintain computational feasibility. However, such approximations either still show limitations in terms of efficiency, when compared to first order methods, or adopt drastic simplifications that make the estimation process less stable.
2.2 Optimization via Kalman Filtering
The idea of using Kalman filtering for optimization has a long history. Early works include scalar Kalman updates for stochastic optimization ismail2018estimation , applications in reinforcement learning shashua2019trust , and connections between the Extended Kalman Filter (EKF) and natural gradient methods ollivier2019extended . Although these approaches demonstrate the versatility of Kalman-based updates, they are often task-specific or limited to small-scale models, and computing full covariance updates remains prohibitively expensive.
More recent research has revisited the Kalman filtering perspective in the context of large-scale neural optimization. Duran-Martin et al. duran2022efficient and Chang et al. chang2023low interpret learning as a Bayesian state estimation problem, where the network parameters are treated as latent states, and the likelihood of the data provides an explicit probabilistic observation model. These methods apply low-rank or subspace-structured versions of the Extended Kalman Filter to maintain tractable posterior updates. Their formulation operates in the full parameter or feature space (i.e., with matrix-valued state and observation models), which implies a computational cost of $\mathcal{O}(n^{2})$ to $\mathcal{O}(n^{3})$ per update, where $n$ is the number of model parameters, even under low-rank approximations. Consequently, they are mainly applicable to some relatively small scale settings.
In contrast, KOALA davtyan2022koala reinterprets gradient-based optimization as a scalar state estimation problem, where the loss function itself serves as an implicit observation. This design removes the need for an explicit probabilistic target and allows Kalman-style uncertainty propagation to be applied directly to general stochastic objectives. The formulation therefore scales linearly with the number of parameters, achieving $\mathcal{O}(n)$ complexity in practice, while remaining compatible with arbitrary differentiable loss functions. This implicit observation model makes KOALA more general and applicable to a wider range of optimization problems than Kalman filters formulated on explicit probabilistic observations. Nevertheless, the original KOALA formulation still involves certain simplifying assumptions that limit its expressiveness and stability in complex, anisotropic optimization landscapes.
For comparison, SLANG mishkin2018slang focuses purely on efficient covariance estimation through the diagonal plus low-rank structure without invoking any Kalman filtering or Bayesian state-update interpretation. Although SLANG achieves improved posterior approximations compared to diagonal methods, its updates are not dynamically informed by temporal prediction-correction as in Kalman filtering, and its computational cost remains at least $\mathcal{O}(n^{2})$.
In general, these methods differ in their dimensionality, observation modeling, and scalability. A gap remains between the expressiveness of structured covariance estimation and the efficiency of scalar Kalman filtering. This motivates the design of KOALA++, which preserves KOALA’s first-order scalability while introducing directionally aware covariance updates for large-scale optimization.
@@ 2.3 Motivation for KOALA++: Towards Structured Covariance Updates
The original KOALA formulation addresses efficiency by interpreting optimization as a scalar Kalman filtering problem, where the loss function acts as an implicit observation. However, its covariance update assumes $P_{k}\approx\sigma^{2}I$ and substitutes $H_{k}^{\top}H_{k}$ with gradient norms, which may reduce some information to a single scalar, limiting its ability to adapt to anisotropic curvature or correlated gradient directions, particularly in deep or noisy regimes.
On the other hand, as discussed in subsection 2.2, better approximations of the covariance matrix can provide richer uncertainty representations, but they are computationally prohibitive. This trade-off leaves a gap between expressiveness and scalability that current methods have yet to bridge.
Thus, KOALA++ is designed to overcome this limitation while retaining the first-order scalability of KOALA. Rather than maintaining or constraining a full covariance matrix, KOALA++ explicitly tracks a directional covariance projection
$v_{k}:=H_{k}P_{k-1},$
which quantifies how uncertainty evolves along current gradient directions. This projection is updated via an approximate recursive least-squares step that implicitly captures curvature anisotropy without requiring explicit matrix storage or inversion. As a result, KOALA++ provides a lightweight mechanism for propagating structured uncertainty, offering an effective balance between expressiveness and computational efficiency.
@@ 3.1 Problem Formulation and Kalman Filtering Notation
Neural network training is fundamentally an optimization problem, where, given a dataset ${\cal D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, the objective of the learning task is to adjust the model’s parameters $\theta\in\mathbb{R}^{n}$ so that the network’s predictions $f(x_{i};\theta)$ minimize the empirical risk
$\min_{\theta}{\cal L}(\theta),\quad\text{with }{\cal L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}l(f(x_{i};\theta),y_{i}),$ (1)
where $l$ is a chosen loss function. However, in modern deep learning applications, $N$ is often large, which makes full-batch optimization computationally infeasible. As a result, practitioners rely on stochastic optimization methods that work on minibatches of data and minimize estimates of the loss
@@
where $B_{k}$ is a set of indices in the $k$-th minibatch and $m=|B_{k}|$ is the cardinality of the minibatch. Although minibatch-based approaches such as SGD robbins1951stochastic and its adaptive variants are effective and scalable, they introduce noise due to the stochastic nature of loss and gradient estimates.
@@
$S_{k}=H_{k}(P_{k-1}+QI)H_{k}^{\top}+R$ innovation covariance, where $H_{k}=\nabla L_{k}^{\top}(\theta_{k-1})$
$K_{k}=(P_{k-1}+QI)H_{k}^{\top}S_{K}^{-1}$ Kalman gain
$P_{k}=(I-K_{k}H_{k})(P_{k-1}+QI)$ posterior of the state update
$\theta_{k}=\theta_{k-1}+K_{k}(L_{k}^{\text{target}}-L_{k}(\theta_{k-1}))$ state estimate update
To mitigate this issue, a recent approach, KOALA davtyan2022koala , casts neural network training as a recursive state estimation problem using Kalman filtering kalman1960new , a principled Bayesian framework for recursively estimating the hidden state of a system from noisy observations. At each iteration $k$, it combines prior estimates with new measurements to update the state $\theta_{k}$ and its associated uncertainty in the form of a covariance matrix $P_{k}$. Importantly, by keeping track of an estimate of the system’s state’s covariance, Kalman filtering is essentially implicitly using second order information about the state variables. The standard Extended Kalman Filter (EKF) generalizes this to non-linear systems using local linearization. In KOALA, model parameters $\theta_{k}$ are treated as latent states with stationary evolution under process noise, and the minibatch loss function $L_{k}$ serves as a noisy observation of a target loss value $L_{k}^{\text{target}}$. In these settings, the EKF equations can be summarized as in Table 1. To remain tractable in high-dimensional settings, KOALA simplifies the EKF update by:
Assuming an identity-scaled covariance matrix $P_{k}\approx\sigma_{k}^{2}I$; Approximating $H_{k}^{\top}H_{k}$ with $\|H_{k}\|^{2}I$; Reducing all matrix operations to scalar computations.
@@
While KOALA demonstrates promising stability and simplicity, its reliance on scalar approximations limits its capacity to capture directional or structured noise in high-dimensional loss landscapes. To address this, we propose KOALA++, which enhances the original framework with an efficient estimation of the directional state covariance while preserving the Kalman-based update algorithm.
@@ 3.2 Kalman Update with $v_{k}$-Based Reparameterization
To overcome the limitations of isotropic covariance assumptions in KOALA, KOALA++ introduces a more expressive formulation that captures directional uncertainty without incurring the full cost of matrix-valued updates. Instead of updating the full posterior covariance $P_{k}\in\mathbb{R}^{n\times n}$, we propagate a low-rank surrogate
@@
which represents the projection of the parameter covariance along the current measurement direction. This allows KOALA++ to maintain a richer understanding of uncertainty and curvature structure while keeping the computational complexity comparable to first-order methods.
@@
Notice that, as suggested in davtyan2022koala , the learning rate of the update can be controlled by selecting a specific $L_{k}^{\text{target}}$. For example, $L_{k}^{\text{target}}=(1-\eta_{k})L_{k}(\theta_{k-1})$ ensures that the update is scaled by $\eta_{k}$.
@@
$=\left(H_{k}-H_{k}\frac{v_{k-1}^{\top}+QH_{k-1}^{\top}}{S_{k-1}}H_{k-1}\right)(P_{k-2}+Q)$ (7)
$=H_{k}P_{k-2}+H_{k}Q-H_{k}\frac{v_{k-1}^{\top}+QH_{k-1}^{\top}}{S_{k-1}}v_{k-1}-H_{k}\frac{v_{k-1}^{\top}+QH_{k-1}^{\top}}{S_{k-1}}H_{k-1}Q.$
For simplicity, we define the recurring term
$\lambda_{k}$ $\doteq H_{k}\frac{v_{k-1}^{\top}+QH_{k-1}^{\top}}{S_{k-1}},$ (8)
which allows us to rewrite the update for $v_{k}$ in a more compact way:
@@
In the last equation, all terms except the first can be efficiently calculated in iteration $k$. The first term cannot be easily evaluated due to the term $P_{k-2}$. The objective is to never store it explicitly to avoid the quadratic storage demand. As a first attempt, one could approximate $H_{k}$ with $H_{k-1}$ and replace the problematic term with $v_{k-1}$. However, we observed experimentally that this does not work well because of the high variability of the minibatch gradients.
As an alternative, we aim to retrospectively approximate the prior covariance matrix $P_{k-2}$ instead. We do this by taking advantage of the relation $v_{k-1}=H_{k-1}P_{k-2}$. However, since this equation is generally underdetermined in high-dimensional settings, we formulate a regularized inverse problem to obtain a unique and tractable solution. Specifically, we adopt the minimum Frobenius-norm formulation
@@
For computational and memory efficiency, we use only one linear constraint, $v_{k-1}=H_{k-1}P_{k-2}$, although more could be used in principle. However, introducing more constraints, e.g., projections of other previous gradients $H_{k-2},H_{k-3}$, etc., would also require memorizing more gradient terms and using more calculations to update $v_{k}$. The use of the minimum Frobenius-norm criterion is to ensure a stable update for $v_{k}$ by limiting the norm of $P_{k-2}$.
In the next subsection, we discuss two variants of our algorithm, depending on the nature of the solution to the optimization problem (10).